# natural language processing; greenwashing detection; corporate climate-related greenwashing; climate pretrained models; climate topic detection; green claim detection; climate QA; climate-related deceptive techniques; enviromental performance prediction copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language processing

### v1

@@ 4.3. Thematic Analysis
Given an input sentence or a paragraph, output a subtopic related to climate change. This is a multiclass classification task that can be either supervised or unsupervised, depending on the availability of labeled data. Alternatively, it can be framed as a clustering task with the goal of discovering latent subtopic structures.
@@
TCFD rec. (Bingler et al., 2022) paragraphs from corporate annual reports TCFD 4 main categories: Metrics and Targets, Risk Management, Strategy, Governance and General (see appendix A for details)
@@
TCFD (Sampson et al., 2022) paragraphs from regulatory and discretionary reports TCFD 11 recommendations: Metrics and Targets (board’s oversight, management’s role), Risk Management (identified risk, impact, resilience), … (see Appendix A for details)
@@
FineBERT’s ESG (Huang et al., 2020) Sentences from 10-K Environmental, Social, Governance, General Environmental - e.g., climate change, natural capital, pollution and waste, and environmental opportunities
@@
ESGBERT’s ESG (Schimanski et al., 2023b) Sentences from reports and corporate news Environment, Social, Governance and None Environmental criteria comprise a company’s energy use, waste management, pollution, […] as well as compliance with governmental regulations. Special areas of interest are climate change and environmental sustainability.
@@
Multilingual ESG classification (Lee and Kim, 2023) Sentences from Korean reports (English and Korean) Environment, Social, Governance, Neutral, and Irrelevant Environmental factors include the reduction of hazardous substances, eco-friendly management, climate change, carbon emission, natural resources, […]
@@
SciDCC (Mishra and Mittal, 2021) News articles (Title, Summary, Body) Environment, Geology, Animals, Ozone Layer, Climate, etc. Category in which the article was published (Automatic Label)
@@
Transaction Ledger (Jain et al., 2023) Transaction ledger entry (description of transaction) Accounting Adjustments, Administration, Advertising, Benefits & Insurance, etc. Standardized commodity classes
@@
ClimateEng (Vaid et al., 2022) Tweets posted during COP25 filtered by keywords (relevant to climate-change) Ocean/Water, Politics, Disaster, Agriculture/Forestry, General Sub-categories of climate-change
@@
ESGBERT’s Nature (Schimanski et al., 2024a) Paragraphs from reports General, Nature, Biodiversity, Forest, Water Multi-label Nature-related topics
@@
ClimaTOPIC (Spokoyny et al., 2023) CDP responses (short texts) Adaptation, Buildings, Climate Hazards, Emissions, Water, etc. Category of the question (Automatic Label)
Table 3. Summary of datasets, labels, scopes, and sources for climate-related text classification tasks.
4.3.1. Classifying Climate-Related Financial Disclosure (TCFD)
@@
The Task Force on Climate-Related Financial Disclosures (TCFD) proposes four main categories of climate-related disclosure (Governance, Strategy, Risk Management, and Metrics and Targets) and 11 recommendations for climate-related disclosures (see Appendix A for details). Finding texts associated to these categories in company communication can help identify gaps and inconsistencies in reporting.
Unsupervised Solutions
@@
Ding et al. (2023) introduced a score to quantify the extent of climate-related content in corporate disclosures, as well as 4 TCFD-category-based similarity scores. The score is based on sentence similarity with sentences from TCFD referential documents. They identified discrepancies that could signal greenwashing by comparing these scores with actual carbon emission data. Auzepy et al. (2023) examined reports from banks endorsing the TCFD recommendations. They proposed a fine-grained analysis of TCFD categories using a zero-shot entailment classification method on sentences.
@@
Although unsupervised methods can be employed for thematic analysis, recent studies have demonstrated the value of curated datasets for categorizing corporate disclosures. For instance, Bingler et al. (2022) leveraged report headings to annotate paragraphs into the four main TCFD categories. Similarly, Sampson et al. (2022) assessed the quality of climate-related disclosures using a dataset of 162k sentences, where each sentence was labeled according to the TCFD recommendations by experts through a question answering (QA) process.
4.3.2. Classifying Environmental, Social, and Governance Disclosure (ESG)
Beyond climate-specific reporting, Environmental, Social, and Governance (ESG) disclosures represent a broader set of non-financial metrics that evaluate a company’s sustainability performance. ESG reporting has become a global standard, encompassing climate-related initiatives and other dimensions of corporate responsibility. The standardized nature of ESG metrics makes them valuable for computational models that aim to detect inconsistencies in corporate claims, such as overstating environmental impacts through greenwashing. However, as noted in Berg et al. (2022)’s aggregated analysis of ESG ratings, discrepancies persist across different evaluation frameworks. These inconsistencies raise the question of whether artificial intelligence can be harnessed to improve the consistency and reliability of ESG assessments.
Unsupervised Solutions
Rouen et al. (2023) proposed an ML algorithm to describe the content of ESG reports. They used industry-topic dictionaries to compute the topic frequency using a TF-IDF-based algorithm. They used that classification to identify selective disclosure. Mehra et al. (2022) experimented with sentence similarity to extract sentences that were the most relevant to environmental factors, and (Bronzini et al., 2023) selected sentences related to ESG using INSTRUCTOR-xl (Su et al., 2023). They both subsequently used the extracted sentences to predict ESG scores.
@@ 4.3.2. Classifying Environmental, Social, and Governance Disclosure (ESG)
Multiple works (Huang et al., 2020; Schimanski et al., 2023b; Lee and Kim, 2023) proposed datasets for topic classification with the labels: Environment, Social, Governance and General. Huang et al. (2020) focused on reports while Schimanski et al. (2023b) added corporate news. Schimanski et al. (2023b) introduced a larger dataset for pre-training domain-specific models and an evaluation dataset of 2,000 sentences. While most ESG-related reports are typically available in English, some local companies provide them only in their native language. To partially address this, Lee and Kim (2023) proposed a multilingual dataset constructed from Korean corporate reports.
4.3.3. Other Topics
@@
Unsupervised Solutions
Kang and Kim (2022) analyzed the content of sustainable reports with thematic analysis. It was conducted using sentence similarity between the reports and content from the SDGs website. Yu et al. (2024) examined the themes covered in the climate-related sections of bank reports by employing a dual approach: using hand-selected keywords to target anticipated topics and applying clustering techniques to group climate-related sentences and uncover underlying themes. Brié et al. (2022) leveraged ClimateBERT to extract climate-related paragraphs from corporate documents. They then applied a Structural Topic Model (STM) (Margaret E. Roberts and Airoldi, 2016) to identify thematic clusters within the data. This STM-based approach was also used by Fortes and Le Guenedal (2020) to identify influential topics for EUR/USD exchange rate. Their findings highlighted a significant discrepancy between the European Central Bank (ECB) and the Federal Reserve (FED), with the ECB attributing greater importance to climate change in the context of financial stability than the FED. Such thematic analysis can be used to identify selective disclosure.
@@
Spokoyny et al. (2023) introduced ClimaTOPIC, a topic classification dataset derived from responses to the CDP questionnaires, organized into 12 categories that align with the thematic structure of the CDP questions. This can be used to identify sections relevant to a particular CDP topic (e.g., Emissions, Adaptation, or Energy), enabling the extraction of structured information from unstructured documents. Schimanski et al. (2024a) presents a dataset focused on nature-related topics (such as Water, Forest, and Biodiversity) extracted from annual reports, sustainability reports, and earnings call transcripts (ESGBERT Nature). While less explicitly linked to climate change, these topics often reflect the consequences of environmental degradation associated with climate issues. While NLP is often used to analyze corporate report texts, it can also be used to analyze tables. For example, Jain et al. (2023) proposes estimating Scope 3 emissions  [^3] Scope 3 emissions refer to all indirect greenhouse gas (GHG) emissions that occur in a company’s value chain, excluding emissions from the company’s own operations (Scope 1) and its purchased electricity (Scope 2). based on corporate transactions (see Section 4.10 for details). The first step involves grouping expenses into standardized commodity classes, for which they introduced a dataset of 8K examples of corporate expenses categorized accordingly. As previously discussed, social media and news articles offer a broader perspective for cross-referencing corporate communications. For example, Vaid et al. (2022) introduced ClimateEng, a climate-related topic classification dataset composed of tweets grouped into five categories: Disaster, Ocean/Water, Agriculture/Forestry, Politics, and General. This dataset can be used to detect greenwashing by comparing the public’s concerns with the narrative presented in corporate sustainability reports. Similarly, Mishra and Mittal (2021) published the Science Daily Climate Change dataset (SciDCC), consisting of 11,000 news articles grouped into 20 climate-related categories. This dataset offers another avenue for identifying greenwashing, as companies’ communications can be cross-referenced with external news coverage, potentially highlighting contradictions or omissions in corporate reports.
4.3.4. Models and Conclusion
Supervised Solutions
@@ 4.3.4. Models and Conclusion
The most common solution proposed for these tasks is fine-tuning a Transformer model (Huang et al., 2020; Schimanski et al., 2023b; Lee and Kim, 2023; Bingler et al., 2022; Sampson et al., 2022; Vaid et al., 2022; Spokoyny et al., 2023; Schimanski et al., 2024a; Jain et al., 2023). To contextualize the performance of Transformer models, some studies also report performances of classical baselines (Huang et al., 2020; Spokoyny et al., 2023; Schimanski et al., 2024a; Jain et al., 2023; Bingler et al., 2022; Sampson et al., 2022). Bingler et al. (2022) proposed a custom approach combining logistic regression with features from a fine-tuned language model, while Sampson et al. (2022) also evaluated clustering techniques and stacked models. The performance of the best-performing model is systematically high (above 80%), except on SciDCC and ClimaTOPIC. Classical approaches (SVM, NB) also reach good performances, under-performing by less than 20% on all datasets (except on Bingler et al. (2022)’s dataset). The TF-IDF baseline even outperformed fine-tuned Transformers on Sampson et al. (2022)’s dataset. This shows that those topics have distinguishable vocabularies.
@@
The performance gap observed in Bingler et al. (2022) may be attributed to using paragraph-level inputs rather than sentence-level inputs aggregated for classification, as fine-tuned models also achieve around 20% precision at the paragraph level. The lower performances on SciDCC(Mishra and Mittal, 2021) and ClimaTOPIC(Spokoyny et al., 2023) are likely caused by the automation of the labeling process. The labels are therefore not designed to be labels. In SciDCC there are labels that are highly similar (e.g. Endangered Animals and Extinctions), or labels that include other labels (e.g. Environment, Climate, Pollution). As they are categories from a journal, the categories changed in time. For ClimaTOPIC, the labels are question categories, which are designed to group questions not to precisely identify them. Therefore a question about ”the emission of a building” might fit in either Emissions or Building, yet it is assigned only one label.
@@
Topic classification is a well-established area of research in natural language processing (NLP) and has been extensively applied to climate-related topics. Fine-tuned Transformer models proposed in the literature consistently demonstrate near-perfect performance on human-annotated datasets, indicating that the task is largely solved. When reported, classical keyword baselines also perform relatively well, achieving significantly better than random performance and approximately 80% of the performance of fine-tuned models, suggesting the topics have distinct vocabularies. Future research could investigate topics that are lexically similar but distinct in their subject or focus. Topic detection serves as a valuable tool for structuring documents and has been utilized to analyze the extent to which companies address specific topics, enabling the detection of selective disclosure (Bingler et al., 2023, 2022). However, it is critical to ensure that labels are well-defined and data is accurately annotated. Lower performance observed on automatically generated labels may reflect challenges in predictability, potentially due to difficulties caused by the automatic annotation. These works might be revisited to conduct human annotations.
4.4. In-depth Disclosure: Climate Risk Classification
@@
Ask BERT’s Climate Risk (Kölbel et al., 2021) Sentences from TCFD’s example reports and non-climate-related sentences Risk type: Transition risk, physical risk, and general risk (no guidelines)
Climate Risk (Friederich et al., 2021) Paragraphs from European companies annual reports Risk type:Acute, Chronic, Policy & legal, Tech & Market, Reputational, and Negative (no guidelines)
ClimateBERT’s Sentiment (Bingler et al., 2023) Paragraphs from companies’ annual reports Risk or threat that negatively impacts an entity of interest (negative sentiment); or Opportunity arising due to climate change (positive sentiment); Neutral otherwise.
Sentiment Analysis (Xiang and Fujii, 2023) paragraphs from academic texts on climate change and health published between 2013 and 2020 Risk (negative) if it discusses climate change causing public health issues, serious consequences, or worsening trends, greenwashing. Opportunity (positive): highlights potential benefits, positive actions, or research addressing gaps. Neutral otherwise.
Table 4. Summary of datasets, labels, scopes, and sources for climate-related Risk classification.
Unsupervised Solutions
Similarly to climate-related detection, climate risk classification has been tackled with keyword-based approaches. Li et al. (2020), Kheradmand et al. (2021) and Chou et al. (2023) proposed using dictionaries of words related to climate risk to identify paragraphs dealing with climate risk. Li et al. (2020) constructed risk measures based on the frequency of the terms in the risk dictionaries. Chou et al. (2023) analyzed the topics mentioned in conjunction with physical and transition risks. Sautner et al. (2023) proposed using one dictionary for risk and one for opportunity classification.
@@
Several studies have introduced specialized annotated datasets for risk/opportunity classification. For instance, Kölbel et al. (2021) created a dataset for climate-related risk classification with three categories: physical risk, transition risk, and general risk, using active learning for annotation. Similarly, Friederich et al. (2021) developed an annotated dataset for risk classification with five labels, covering acute physical risk, chronic physical risk, policy and legal risks, technology and market risks, and reputational transition risks. Bingler et al. (2023) released a dataset as part of the ClimateBERT downstream tasks, focusing on classifying paragraphs from corporate reports into three categories: opportunity, neutral, or risk. Extending beyond corporate disclosures, Xiang and Fujii (2023) compiled a climate-related risk/opportunity classification dataset of academic texts from the Web of Science and Scopus.
Supervised Solutions
The solutions proposed are similar to the other tasks: fine-tuned Transformer models (Kang and Kim, 2022; Kölbel et al., 2021; Friederich et al., 2021; Bingler et al., 2023; Webersinke et al., 2022; Xiang and Fujii, 2023) and keyword-based-features with simple models (Kölbel et al., 2021; Friederich et al., 2021; Bingler et al., 2023) (see Table LABEL:tab:models in the appendix for details). The notable exception is Xiang and Fujii (2023), who evaluated LSTM-based solutions alongside the Transformers models. Most studies (Kang and Kim, 2022; Kölbel et al., 2021; Bingler et al., 2023; Webersinke et al., 2022; Xiang and Fujii, 2023) reported high performances, above 80%. Kölbel et al. (2021) and Bingler et al. (2023) also reported good performances for the keyword-based baselines (72% and 84%). The performance reported by Friederich et al. (2021) on climate risk tasks highlights distinct challenges across subtasks. When tasked with identifying the specific type of risk in sentences already known to be about climate risks, word-based models outperformed fine-tuned Transformers. This suggests that each risk type had a distinct vocabulary. In contrast, when working with a dataset that reflects real-world proportions of risk and non-risk examples—where the data is heavily imbalanced—classical models struggled due to the overlap in vocabulary between general and climate risk-related text.
@@
It is important to distinguish risk/opportunity classification from sentiment analysis. While risk/opportunity classification is inspired by sentiment analysis (Bingler et al., 2023), they play a complementary role. A statement such as “With proactive climate risk management, we are ready to tackle extreme weather disruptions, ensuring resilience” is classified as positive using a sentiment analysis model (Pérez et al., 2021), and as mentioning a risk using a risk/opportunity classification model (Bingler et al., 2023). Sentiment analysis focuses on the form of the statement, while risk is about the content. As the literature on sentiment analysis (Wankhade et al., 2022) is broad, the analysis of the tone of the statement can be done using existing models (Polignano et al., 2022; Kang and Kim, 2022) complementing approaches on risk classification.
@@
The performance of fine-tuned models might indicate that the task of identifying texts talking about risks is solved. However, the work by Friederich et al. (2021) shows limitations when experimenting with heavily imbalanced datasets, which actually correspond to real-world settings. Future research could focus on such skewed settings.
4.5. Green Claim Detection
Identifying greenwashing involves more than just finding text related to climate change, as a company might simply state factual information about the climate without making any commitments on it. Hence, we now focus on detecting green claims, i.e., claims that a product, service, or corporate practice either contributes positively to environmental sustainability or is less harmful to the environment than alternatives (Stammbach et al., 2023; Woloszyn et al., 2021). The European Commission calls these claims “environmental claims”:

### v2

@@ 4.3 Thematic Analysis
Given an input sentence or paragraph, output a subtopic related to climate change. This is a multiclass classification task that can be supervised or unsupervised, depending on the availability of labeled data. Alternatively, it can be framed as a clustering task with the goal of discovering latent subtopic structures.
Table 5: Summary of datasets, labels, scopes, and sources for climate-related text classification tasks.
@@
TCFD rec. bingler_cheap_2021 Paragraphs from corporate annual reports TCFD 4 main categories: Metrics and Targets, Risk Management, Strategy, Governance and General (see appendix A for details)
@@
TCFD sampson_tcfd-nlp_nodate Paragraphs from regulatory and discretionary reports TCFD 11 recommendations: Metrics and Targets (board’s oversight, management’s role), Risk Management (identified risk, impact, resilience), … (see Appendix A for details)
@@
FineBERT’s ESG huangFinBERTLargeLanguage2020 Sentences from 10-K Environmental, Social, Governance, General Environmental - e.g., climate change, natural capital, pollution and waste, and environmental opportunities
@@
ESGBERT’s ESG schimanski_bridging_2023 Sentences from reports and corporate news Environment, Social, Governance and None Environmental criteria comprise a company’s energy use, waste management, pollution, […] as well as compliance with governmental regulations. Special areas of interest are climate change and environmental sustainability.
@@
Multilingual ESG classification LEE2023119726 Sentences from Korean reports (English and Korean) Environment, Social, Governance, Neutral, and Irrelevant Environmental factors include the reduction of hazardous substances, eco-friendly management, climate change, carbon emission, natural resources, […]
@@
SciDCC mishra2021neuralnere News articles (Title, Summary, Body) Environment, Geology, Animals, Ozone Layer, Climate, etc. Category in which the article was published (Automatic Label)
@@
Transaction Ledger jain_supply_2023 Transaction ledger entry (description of transaction) Accounting Adjustments, Administration, Advertising, Benefits & Insurance, etc. Standardized commodity classes
@@
ClimateEng vaid-etal-2022-towards Tweets posted during COP25 filtered by keywords (relevant to climate-change) Ocean/Water, Politics, Disaster, Agriculture/Forestry, General Sub-categories of climate-change
@@
ESGBERT’s Nature Schimanski2024nature Paragraphs from reports General, Nature, Biodiversity, Forest, Water Multi-label Nature-related topics
@@
ClimaTOPIC spokoyny2023answering CDP responses (short texts) Adaptation, Buildings, Climate Hazards, Emissions, Water, etc. Category of the question (Automatic Label)
4.3.1 Classifying Climate-Related Financial Disclosure (TCFD)
@@
The Task Force on Climate-Related Financial Disclosures (TCFD) proposes four main categories of climate-related disclosure (Governance, Strategy, Risk Management, and Metrics and Targets) and 11 recommendations for climate-related disclosures.444see Appendix A for details. Finding texts associated to these categories in company communication can help identify gaps and inconsistencies in reporting.
Annotation-free
@@
dingCarbonEmissionsTCFD2023 introduced a score to quantify the extent of climate-related content in corporate disclosures, as well as 4 TCFD-category-based similarity scores. The score is based on sentence similarity with sentences from TCFD referential documents. They identified discrepancies that could signal greenwashing by comparing these scores with actual carbon emission data. auzepy_evaluating_2023 examined reports from banks endorsing the TCFD recommendations. They proposed a fine-grained analysis of TCFD categories using a zero-shot entailment classification method on sentences.
@@
Although unsupervised methods can be employed for thematic analysis, recent studies have demonstrated the value of curated datasets for categorizing corporate disclosures. For instance, bingler_cheap_2021 leveraged report headings to annotate paragraphs into the four main TCFD categories. Similarly, sampson_tcfd-nlp_nodate assessed the quality of climate-related disclosures using a dataset of 162k sentences, where each sentence was labeled according to the TCFD recommendations by experts through a question answering (QA) process.
4.3.2 Classifying Environmental, Social, and Governance Disclosure (ESG)
Beyond climate-specific reporting, Environmental, Social, and Governance (ESG) disclosures represent a broader set of non-financial metrics that evaluate a company’s sustainability performance. ESG reporting has become a global standard, encompassing climate-related initiatives and other dimensions of corporate responsibility. The standardized nature of ESG metrics makes them valuable for computational models that aim to detect inconsistencies in corporate claims, such as overstating environmental impacts through greenwashing. However, as noted in berg2022aggregate’s aggregated analysis of ESG ratings, discrepancies persist across different evaluation frameworks. These inconsistencies raise the question of whether artificial intelligence can be harnessed to improve the consistency and reliability of ESG assessments.
Annotation-free
rouenEvolutionESGReports2023 proposed an ML algorithm to describe the content of ESG reports. They used industry-topic dictionaries to compute the topic frequency using a TF-IDF-based algorithm. They used that classification to identify selective disclosure. Mehra_2022 experimented with sentence similarity to extract sentences that were the most relevant to environmental factors, and bronzini_glitter_2023 selected sentences related to ESG using INSTRUCTOR-xl [su2023embeddertaskinstructionfinetunedtext]. They both subsequently used the extracted sentences to predict ESG scores.
@@ 4.3.2 Classifying Environmental, Social, and Governance Disclosure (ESG)
Multiple works huangFinBERTLargeLanguage2020, schimanski_bridging_2023, LEE2023119726 proposed datasets for topic classification with the labels: Environment, Social, Governance and General. huangFinBERTLargeLanguage2020 focused on reports while schimanski_bridging_2023 added corporate news. schimanski_bridging_2023 introduced a larger dataset for pre-training domain-specific models and an evaluation dataset of 2,000 sentences. While most ESG-related reports are typically available in English, some local companies provide them only in their native language. To partially address this, LEE2023119726 proposed a multilingual dataset constructed from Korean corporate reports.
4.3.3 Other Topics
@@
Annotation-free
hyewon_kang_analyzing_2022 analyzed the content of sustainable reports with thematic analysis. It was conducted using sentence similarity between the reports and content from the SDGs website. yu_climatebug_2024 examined the themes covered in the climate-related sections of bank reports by employing a dual approach: using hand-selected keywords to target anticipated topics and applying clustering techniques to group climate-related sentences and uncover underlying themes. bjarne_brie_mandatory_2022 leveraged ClimateBERT to extract climate-related paragraphs from corporate documents. They then applied a Structural Topic Model (STM) [STM] to identify thematic clusters within the data. This STM-based approach was also used by fortes2020tracking to identify influential topics for EUR/USD exchange rate. Their findings highlighted a significant discrepancy between the European Central Bank (ECB) and the Federal Reserve (FED), with the ECB attributing greater importance to climate change in the context of financial stability than the FED. Such thematic analysis can be used to identify selective disclosure.
@@
spokoyny2023answering introduced ClimaTOPIC, a topic classification dataset derived from responses to the CDP questionnaires, organized into 12 categories that align with the thematic structure of the CDP questions. This can be used to identify sections relevant to a particular CDP topic (e.g., Emissions, Adaptation, or Energy), enabling the extraction of structured information from unstructured documents. Schimanski2024nature presents a dataset focused on nature-related topics (such as Water, Forest, and Biodiversity) extracted from annual reports, sustainability reports, and earnings call transcripts (ESGBERT Nature). While less explicitly linked to climate change, these topics often reflect the consequences of environmental degradation associated with climate issues. While NLP is often used to analyze corporate report texts, it can also be used to analyze tables. For example, jain_supply_2023 proposes estimating Scope 3 emissions  [^5] Scope 3 emissions refer to all indirect greenhouse gas (GHG) emissions that occur in a company’s value chain, excluding emissions from the company’s own operations (Scope 1) and its purchased electricity (Scope 2). based on corporate transactions (see Section 4.10 for details). The first step involves grouping expenses into standardized commodity classes, for which they introduced a dataset of 8K examples of corporate expenses categorized accordingly. As previously discussed, social media and news articles offer a broader perspective for cross-referencing corporate communications. For example, vaid-etal-2022-towards introduced ClimateEng, a climate-related topic classification dataset composed of tweets grouped into five categories: Disaster, Ocean/Water, Agriculture/Forestry, Politics, and General. This dataset can be used to detect greenwashing by comparing the public’s concerns with the narrative presented in corporate sustainability reports. Similarly, mishra2021neuralnere published the Science Daily Climate Change dataset (SciDCC), consisting of 11,000 news articles grouped into 20 climate-related categories. This dataset offers another avenue for identifying greenwashing, as companies’ communications can be cross-referenced with external news coverage, potentially highlighting contradictions or omissions in corporate reports.
4.3.4 Models and Conclusion
Models
@@ 4.3.4 Models and Conclusion
The most common solution proposed for these tasks is fine-tuning a Transformer model [huangFinBERTLargeLanguage2020, schimanski_bridging_2023, LEE2023119726, bingler_cheap_2021, sampson_tcfd-nlp_nodate, vaid-etal-2022-towards, spokoyny2023answering, Schimanski2024nature, jain_supply_2023]. To contextualize the performance of Transformer models, some studies also report performances of classical baselines [huangFinBERTLargeLanguage2020, spokoyny2023answering, Schimanski2024nature, jain_supply_2023, bingler_cheap_2021, sampson_tcfd-nlp_nodate]. bingler_cheap_2021 proposed a custom approach combining logistic regression with features from a fine-tuned language model, while sampson_tcfd-nlp_nodate also evaluated clustering techniques and stacked models. The performance of the best-performing model is systematically high (above 80%), except on SciDCC and ClimaTOPIC. Classical approaches (SVM, NB) also reach good performances, under-performing by less than 20% on all datasets (except on bingler_cheap_2021’s dataset). The TF-IDF baseline even outperformed fine-tuned Transformers on sampson_tcfd-nlp_nodate’s dataset. This shows that those topics have distinguishable vocabularies.
@@
The performance gap observed in bingler_cheap_2021 may be attributed to using paragraph-level inputs rather than sentence-level inputs aggregated for classification, as fine-tuned models also achieve around 20% precision at the paragraph level. The lower performances on SciDCCmishra2021neuralnere and ClimaTOPICspokoyny2023answering are likely caused by the automation of the labeling process. The labels are therefore not designed to be labels. In SciDCC there are labels that are highly similar (e.g. Endangered Animals and Extinctions), or labels that include other labels (e.g. Environment, Climate, Pollution). As they are categories from a journal, the categories changed in time. For ClimaTOPIC, the labels are question categories, which are designed to group questions not to precisely identify them. Therefore, a question about ”the emission of a building” might fit in either Emissions or Building, yet it is assigned only one label.
@@
Topic classification is a well-established area of research in natural language processing (NLP) and has been extensively applied to climate-related topics. Fine-tuned Transformer models proposed in the literature consistently demonstrate near-perfect performance on human-annotated datasets, indicating that the task is largely solved. When reported, classical keyword baselines also perform relatively well, achieving significantly better than random performance and approximately 80% of the performance of fine-tuned models, suggesting the topics have distinct vocabularies. Future research could investigate topics that are lexically similar but distinct in their subject or focus. Topic detection serves as a valuable tool for structuring documents and has been utilized to analyze the extent to which companies address specific topics, enabling the detection of selective disclosure bingler2023cheaptalkspecificitysentiment, bingler_cheap_2021. However, it is critical to ensure that labels are well-defined and data is accurately annotated. Lower performance observed on automatically generated labels may reflect challenges in predictability, potentially due to difficulties caused by the automatic annotation. These works might be revisited to conduct human annotations.
4.4 In-depth Disclosure: Climate Risk Classification
@@
Table 6: Summary of datasets, labels, scopes, and sources for climate-related Risk classification.
@@
Ask BERT’s Climate Risk kolbel_ask_2021 Sentences from TCFD’s example reports and non-climate-related sentences Risk type: Transition risk, physical risk, and general risk (no guidelines)
Climate Risk Friederich_climate_risk_disclosure Paragraphs from European companies annual reports Risk type:Acute, Chronic, Policy & legal, Tech & Market, Reputational, and Negative (no guidelines)
ClimateBERT’s Sentiment bingler2023cheaptalkspecificitysentiment Paragraphs from companies’ annual reports Risk or threat that negatively impacts an entity of interest (negative sentiment); or Opportunity arising due to climate change (positive sentiment); Neutral otherwise.
Sentiment Analysis xiang_dare_2023 Paragraphs from academic texts on climate change and health published between 2013 and 2020 Risk (negative) if it discusses climate change causing public health issues, serious consequences, or worsening trends, greenwashing. Opportunity (positive): highlights potential benefits, positive actions, or research addressing gaps. Neutral otherwise.
Annotation-free
Similarly to climate-related detection, climate risk classification has been tackled with keyword-based approaches. liCorporateClimateRisk2020, kheradmand2021a and chou_ESG proposed using dictionaries of words related to climate risk to identify paragraphs dealing with climate risk. liCorporateClimateRisk2020 constructed risk measures based on the frequency of the terms in the risk dictionaries. chou_ESG analyzed the topics mentioned in conjunction with physical and transition risks. SAUTNER_cliamte_change_exp proposed using one dictionary for risk and one for opportunity classification.
@@
Several studies have introduced specialized annotated datasets for risk/opportunity classification. For instance, kolbel_ask_2021 created a dataset for climate-related risk classification with three categories: physical risk, transition risk, and general risk, using active learning for annotation. Similarly, Friederich_climate_risk_disclosure developed an annotated dataset for risk classification with five labels, covering acute physical risk, chronic physical risk, policy and legal risks, technology and market risks, and reputational transition risks. bingler2023cheaptalkspecificitysentiment released a dataset as part of the ClimateBERT downstream tasks, focusing on classifying paragraphs from corporate reports into three categories: opportunity, neutral, or risk. Extending beyond corporate disclosures, xiang_dare_2023 compiled a climate-related risk/opportunity classification dataset of academic texts from the Web of Science and Scopus.
Models
The solutions proposed are similar to the other tasks: fine-tuned Transformer models hyewon_kang_analyzing_2022, kolbel_ask_2021, Friederich_climate_risk_disclosure, bingler2023cheaptalkspecificitysentiment, nicolas_webersinke_climatebert_2021, xiang_dare_2023 and keyword-based-features with simple models kolbel_ask_2021, Friederich_climate_risk_disclosure, bingler2023cheaptalkspecificitysentiment. The notable exception is xiang_dare_2023, who evaluated LSTM-based solutions alongside the Transformers models. Most studies [hyewon_kang_analyzing_2022, kolbel_ask_2021, bingler2023cheaptalkspecificitysentiment, nicolas_webersinke_climatebert_2021, xiang_dare_2023] reported high performances, above 80%. kolbel_ask_2021 and bingler2023cheaptalkspecificitysentiment also reported good performances for the keyword-based baselines (72% and 84%). The performance reported by Friederich_climate_risk_disclosure on climate risk tasks highlights distinct challenges across subtasks. When tasked with identifying the specific type of risk in sentences already known to be about climate risks, word-based models outperformed fine-tuned Transformers. This suggests that each risk type had a distinct vocabulary. In contrast, when working with a dataset that reflects real-world proportions of risk and non-risk examples—where the data is heavily imbalanced—classical models struggled due to the overlap in vocabulary between general and climate risk-related text.
@@
It is important to distinguish risk/opportunity classification from sentiment analysis. While risk/opportunity classification is inspired by sentiment analysis bingler2023cheaptalkspecificitysentiment, they play a complementary role. A statement such as “With proactive climate risk management, we are ready to tackle extreme weather disruptions, ensuring resilience” is classified as positive using a sentiment analysis model perez2021pysentimiento, and as mentioning a risk using a risk/opportunity classification model bingler2023cheaptalkspecificitysentiment. Sentiment analysis focuses on the form of the statement, while risk is about the content. As the literature on sentiment analysis Wankhade2022 is broad, the analysis of the tone of the statement can be done using existing models marco_polignano_nlp_2022, hyewon_kang_analyzing_2022 complementing approaches on risk classification.
@@
The performance of fine-tuned models might indicate that the task of identifying texts talking about risks is solved. However, the work by Friederich_climate_risk_disclosure shows limitations when experimenting with heavily imbalanced datasets, which actually correspond to real-world settings. Future research could focus on such skewed settings.
4.5 Green Claim Detection
Identifying greenwashing involves more than just finding text related to climate change, as a company might simply state factual information about the climate without making any commitments on it. Hence, we now focus on detecting green claims, i.e., claims that a product, service, or corporate practice either contributes positively to environmental sustainability or is less harmful to the environment than alternatives stammbach_environmental_2023, vinicius_woloszyn_towards_2021. The European Commission calls these claims “environmental claims”:
Table 7: Label definitions of the datasets on green claims detection.